Three. Numbers are in A.P. the product of first and last is 216 and 4 times the second together with twice the first is 84. Find the. NUmbers
Answer by Edwin McCravy(20065)

Three. Numbers are in A.P.
First term = a
Second term = a+d
Third term = a+2d

the product of first and last is 216
a(a+2d) = 216

4 times the second together with twice the first is 84
4(a+d) + 2a = 84

So we have this system of two equations in two unknowns:

    a(a+2d) = 216
  4(a+d)+2a = 84

Simplify the 2nd equation:

  4a+4d+2a = 84
     6a+4d = 84

Divide through by 2

     3a+2d = 42

Solve for 2d
     
        2d = 42-3a

Substitute 42-3a for 2d in
  
   a(a+2d) = 216
a(a+42-3a) = 216
 a(-2a+42) = 216
  -2aČ+42a = 216

Get 0 on the right

-2aČ+42a-216 = 0

Divide through by -2

  aČ-21a+108 = 0
 (a-12)(a-9) = 0

a-12=0;   a-9=0
   a=12;    a=9

Substitute a=12 in
   
         2d = 42-3a
         2d = 42-3(12)
         2d = 42-36
         2d = 6
          d = 3

 That gives the solution

First term = a = 12
Second term = a+d = 12+3 = 15
Third term = a+2d = 12+2(3) = 12+6 = 18

One solution:  12, 15, 18

Substitute a=9 in
   
         2d = 42-3a
         2d = 42-3(9)
         2d = 42-27
         2d = 15
          d = 15/2 = 7.5

 That gives the solution

First term = a = 9
Second term = a+d = 9+7.5 = 16.5
Third term = a+2d = 9+2(7.5) = 9+15 = 24

Other solution:  9, 16.5, 24

Edwin